Assume that in 18 years you would like to have a savings account = $125,000. If you can invest your savings today at 5%, what amount must you invest today to reach your goal?

Assume that you have decided to save $5,000 per year for the next 25 years. If you can earn 9% interest per year on your saving, what will be the value of your wealth in 25 years?
Answer #1

Amount required in 18 years = 125000. Interest rate = 5%. Amount needed to be invested today = PV = FV(1 +

r)^-n which gives PV = 125000 x (1 + 5%)^-18 = $51,940.

Value of wealth in 25 years = 5000 (F/A, 9%, 25) = 5000 x 84.7009 = $423,504.50